Misu sheet,owner of the bedspread shop, knows his customer will pay no more than $150 for a comforter. Misu sheet wants to advertise the comforter as percent markup on cost.
What is the equivalent rate of percent markup on cost compared to the 40% markup on selling price? ( round your answer to the nearest hundreth percent)

Answer by Theo(13342)

if i understand you correctly, he wants to sell the comforter for $150 which would be 40% off the retail price and he also wants to make 40% profit over the cost.

if the discount price = 150 and that price needs to be 40% above the cost, then the cost would be determined as follows:

x = the cost.
x + .40 * x = the discount price = 150
that gets you 1.4 * x = 150
solve for x to get x = 150 / 1.4 = 107.1428571
the cost would be 107.1428571.

y = the retail price.
y - .40 * y = the discount price = 150
that gets you .6 * y = 150
solve for y to get y = 250
the retail price would be 250.

the cost is equal to 107.1428571
the retail price is equal to 250.
the discount price is 150.

the percent markup of the retail price over the cost would be equal to (250 - 107.1428571) / 107.1428571 = 1.33333..... = 133.3333333%

here's how it works.

the cost is 107.1428571.
the markup on the retail price over the cost is 133.3333333%
that makes the retail price equal to 107.1428571 + 1.3333333 * 107.1428571 = 250.

he now wants so sell the comforter at a discount of 40% over the retail price.

250 - .4 * 250 = 150.

he advertises the retail price of the comforter at 250.
he advertises 40% off the retail price which makes the discount price equal to 150.
the discount price of 150 is 40% above the cost of 107.1428571.

your solution appears to be:
the percent markup of the retail price over the cost is 133.33%.
the percent markup of the discount price over the cost is 40%